Consider the initial-value problem y + x(y)2 = 0, y(1) = 4, y(1) = 2

1) Graph the solution curve from [0,5] with step .1

2) Find the explicit solution of the IVP

3) Find the interval of definition
Answer #1

2 +1 2 ( x 2 solving equation by reduction of order. y+x(y)2 =0 , y(1) =4, y(t) = 2. Let y=w wt aw?=0 - du le = -xw2 - s

!! PLEASE RATE WITH THUMBS UP !!